The 2014-15 Brisbane City Council budget included 179 million dollars for public parks. If only 42 percent of Brisbanes 2.1 mllion residents value and use public parks, how much do these residents need to individually value the parks for this to be an efficient use of funds? Answer to the nearest whole dollar per resident (with no decimal points, spaces, S signs, or commas in your answer).
Answer #1

Budget = 179 million

Total residents = 2.1 million

Number of residents value the park = 42% (2.1 million) = 0.42*2.1 million = 0.882 million

Individual value the park for efficient use of funds = budget/ Number of residents value the park

Dollar per resident = 179/0.882

Dollar per resident = 202.95 dollars/ resident

Answer= 203 approximately